# AutoElicit: Using Large Language Models for Expert Prior Elicitation in Predictive Modelling

### v4

@@ 7 Conclusion
Adding expert guidance into the task descriptions enables further improvements to predictive accuracy, as the LLM updates elicited distributions as we would expect. AutoElicit can therefore also be used to simplify the typical expert prior elicitation process, by only requiring experts to provide insights as natural language. We find that prior elicitation with AutoElicit is an improved alternative to in-context learning in our setting, as we can inspect priors, confidently perform Bayesian inference, and often achieve reduced error.
On our private dataset we find a large and quantifiable performance improvement when predicting cases of UTI. Within our study on dementia care the resulting model created with LLM-elicited priors can provide better predictions of adverse health conditions, reliable uncertainty estimates, and remain interpretable and trusted by clinicians.
@@
AC: Original idea conception, software, analysis of results, writing, editing, reviewing, funding acquisition. RGK: Original idea conception, supervision, analysis of results, writing, editing, reviewing, funding acquisition. PB: Supervision, analysis of results, reviewing, funding acquisition.
@@ Acknowledgments
This study is funded by the UK Dementia Research Institute (UKDRI) Care Research and Technology Centre funded by the Medical Research Council (MRC), Alzheimer’s Research UK, Alzheimer’s Society (grant number: UKDRI–7002), and the UKRI Engineering and Physical Sciences Research Council (EPSRC) PROTECT Project (grant number: EP/W031892/1). Infrastructure support for this research was provided by the NIHR Imperial Biomedical Research Centre (BRC) and the UKRI Medical Research Council (MRC). Further funding was provided by the Turing scheme, UK, and resources used in preparing this research were provided, in part, by the Province of Ontario, the Government of Canada through CIFAR, and companies sponsoring the Vector Institute. Additionally, RGK acknowledges support from a Canada CIFAR AI Chair. The funders were not involved in the study design, data collection, data analysis, or manuscript writing.
@@ A.1 Availability of code
Within the supplementary materials, we provide code for generating all figures and results, as well as a pipeline for using our method on additional datasets.
The open-source code makes it simple to perform expert prior elicitation using LLMs: {minted}[mathescape, numbersep=5pt, frame=lines, framesep=5mm, ]python # import the elicitation function from llm_elicited_priors.gpt import ( get_llm_elicitation_for_dataset )
@@
This code is implemented in python (v3.11.9), using pip (v24.0). For our experiments, we use the packages:
@@ A.2.1 Using LLMs to elicit expert priors
Choi et al. (2022) explores the use of an LLM to perform feature selection, causal discovery, and to modify the reward function in a reinforcement learning task. In the first case, they observe that when listing features from two datasets, GPT-3 can correctly identify the features that correspond to a single specified predictive task. Additionally, when applied to causal discovery, the language model is often able to predict the directionality of causal relationships in the Tüebingen Cause-Effect Pairs dataset (Mooij et al., 2016). This inspires us to explore the possibility of using LLMs for eliciting prior parameter distributions over linear models.
Gouk & Gao (2024) use LLMs to generate synthetic data to elicit priors on public datasets. A logistic regression model is then pre-trained using this, defining its prior. However, this work does not discuss the memorisation of public datasets by LLMs, which is shown to be substantial Bordt et al. (2024), supported by the results in Appendix A.8.
Further, Selby et al. (2024) elicit parameters of a beta distribution to define a prior for predictive tasks where results from human experts are available. The priors from four LLMs were visually compared to those elicited from human experts using the SHELF (Oakley & O’Hagan, 2019; Gosling, 2018) method, which demonstrated significant differences between experts and LLMs. However, they show, by considering prior predictive distributions, that GPT-4 (OpenAI et al., 2024) can provide parameter distributions that allow for zero-shot performance that compares to models trained on tens of data points. In our work, we extend this by using the LLM to elicit multiple Gaussian distributions as priors. We combine these to create a mixture model for a more general distribution over our parameters and test their utility for linear predictive models.
In addition, Requeima et al. (2024) prompt language models with task descriptions and collected data points to generate posterior predictive distributions for regression tasks, which they find outperform Gaussian processes in some tasks. This is done by repeatedly sampling a language model at a new feature value, given the previous input and output observations, and a given text description that contains prior information from a human expert.
@@ A.2.2 In-context learning
Learning from demonstrations in-context has proven to be effective for several natural language tasks, demonstrating the surprising ability of language models to perform outside of their intended use (Brown et al., 2020; Dong et al., 2022; Hegselmann et al., 2023; Chowdhery et al., 2023; Touvron et al., 2023). In Brown et al. (2020), the authors study GPT-3’s strong ability on natural language tasks after seeing a few examples, one example, or none at all. This work inspired others to study the extent to which language models can learn from demonstrations given as text alongside a task description. Further, Xie et al. (2021) study the mechanics of in-context learning and describes it from a Bayesian Inference perspective. They find that in-context learning succeeds when language models can infer a shared concept across individual examples in a prompt. Min et al. (2022) argue that in-context learning allows the language model to observe the label space, the distribution of input text, and the task format, as they find that swapping the real labels in demonstrations for random labels does not significantly reduce accuracy. Most of the existing work focuses on natural language tasks Dong et al. (2022), leaving numerical problems less explored and a focus of our work.
In Hollmann et al. (2022), the authors train a transformer model to solve synthetically generated classification tasks from a prior over tabular datasets. Given a new task and demonstrations, this transformer performs in-context learning to make predictions using inference steps only. Furthermore, Akyürek et al. (2022) and Von Oswald et al. (2023) discuss the types of learning algorithms that language models (and more generally, transformers) can implement in-context. Focusing on linear models, Akyürek et al. (2022) show that, by construction, transformers are capable of performing gradient descent and closed-form ridge regression. Akyürek et al. (2022) also demonstrate that trained in-context learners closely match predictors trained through gradient descent. Further, Von Oswald et al. (2023) show that transformers, by incorporating multilayer perceptrons into their architectures, can also solve non-linear regression tasks, and Dalal & Misra (2024) propose a Bayesian framework to reason about a language model’s internal state and suggest that in-context learning is consistent with Bayesian learning. In our work, we test empirically, through the approximation of the internal model used by a language model for its in-context predictions, whether for a given task, the language model is performing Bayesian inference before and after seeing demonstrations.
@@ A.3 Formal details on extracting the internal prior and posterior
In this section, we provide a description of the method presented in Section 4.2 in more detail.
@@
A.4 Does the language model use the model class we specify?
As part of comparing AutoElicit with in-context learning, we use maximum likelihood sampling to extract the internal parameter distributions (discussed in Section 5.4), which we use to test whether the language model is: (1) providing a similar prior distribution through elicitation to the one it is using in-context, and (2) empirically test whether the language model is approximating Bayesian inference. However, when extracting the internal parameters of the in-context model, we assume that the language model is applying the model class that we specify in the task description.
@@ A.4 Does the language model use the model class we specify?
In Figure 8 we show the true $\hat{y}_{n}$ against the in-context predictions $y_{n}$. Interestingly, even given the simple linear equation we ask the language model to mimic, we find that it struggles to provide good predictions in-context. This suggests that the language model is not always using the model class for in-context predictions that we specify in the task description, indicating that in its current state, in-context learning might be inadequate for predictive tasks with primarily numerical features. This figure also shows that the language model was hesitant to predict negative values and preferred positive predictions – a strange behaviour that clearly restricts its use as an emulator of a linear model.
@@ A.7 Further experimental details
In all experiments, we use OpenAI’s GPT-3.5-turbo (specifically GPT-3.5-turbo-0125), GPT-4-turbo (GPT-4-turbo-2024-04-09) [^2] https://platform.openai.com/docs/api-reference/chat, and DeepSeek-R1-32B-Q4333https://ollama.com/library/deepseek-r1:32b with a temperature of $0.1$ and all other settings as default.
@@ A.7.1 Dataset details
  • A private dataset collected as part of our study on Dementia care (Capstick et al., 2024) [^4] https://www.imperial.ac.uk/uk-dri-care-research-technology/: This dataset contains $10$ features engineered from measurements of in-home activity and physiology. The raw features were measured using passive infrared (PIR) sensors located around the homes of people living with dementia, and a sleep mat positioned in the participants bed. The targets correspond to days with a clinical labelling of positive or negative urinary tract infection (UTI) which were calculated through the analysis of a provided urine sample in consultation with a healthcare professional. For this analysis, we used a subset of the data containing $78$ people and $371$ labelled days of UTI, with $270$ negatives and $101$ positives. When training and testing on this dataset, we centred and scaled the training data and used the same statistics to normalise the test data.
@@
  • The Diabetes dataset (Efron et al., 2004): After preprocessing, this dataset contains $442$ people with Diabetes and some physiological measurements. Along with some demographic information, the task is to predict a quantitative measure of their diabetes disease progression one year after this baseline. This dataset consisted of $10$ features and the targets were regression values with a maximum value of $10.0$, minimum value of $0.0$, and a mean value of $3.96$.
@@ A.7.2 Language model details
  • GPT-3.5-turbo [^5] https://platform.openai.com/docs/models is one of OpenAI’s least expensive models and is described as a model for simple tasks.
  • GPT-4-turbo [^6] https://platform.openai.com/docs/models is one of OpenAI’s more sophisticated models.
  • DeepSeek-R1-32B-Q4777https://ollama.com/library/deepseek-r1:32b, a 32 billion parameter model from DeepSeek based on Qwen 32B and distilled from DeepSeek-R1.
@@ A.7.3 Evaluating LLM-elicited priors
To evaluate the usefulness of the AutoElicit priors (presented in Section 4.1), we started by comparing their posterior performance to an uninformative prior. To do this, we chose the predictive tasks described in Section 5.1: $y=2x_{1}-x_{2}+x_{3}$, Urinary Tract Infection (UTI), Heart Disease, Hypothyroidism, Diabetes, and Breast Cancer and calculated the posterior accuracy or mean squared error on a test set after observing varied numbers of examples. For each of the datasets, we selected $50\%$ of data points to form the test set and used the remaining data points to form a training set. This is done $10$ times to produce $10$ folds of random training and testing splits.
@@
To produce the results in Figure 9, we then sample the posterior predictive distribution over our test set and calculate the accuracy or mean squared error for each posterior parameter sample. To produce Figures 2 and 12, we also calculate the mean accuracy and the mean squared error for each of the test splits before plotting the mean and the $95\%$ confidence interval.
As an example, for the synthetic task one of the task descriptions is as follows:
@@
  • User role: I am a data scientist working with a dataset for a prediction task using feature values to predict a target. I would like to apply your model to predict the target for my samples. My dataset consists of these features: [’feature 0’, ’feature 1’, ’feature 2’]. All the values have been standardized using z-scores. The known relationship is: ’target’ = 2 * ’feature 0’ - 1 * ’feature 1’ + 1 * ’feature 2’. Based on the correlation between each feature and target, whether positive or negative, please guess the mean and standard deviation for a normal distribution prior for each feature. I need this for creating a linear regression model to predict the target. Provide your response as a JSON object with the feature names as keys, each containing a nested dictionary for the mean and standard deviation. A positive mean suggests positive correlation with the outcome, negative for negative correlation, and a smaller standard deviation indicates higher confidence. Only respond with JSON.

### v5

@@ 7 Conclusion
Adding expert guidance to the task descriptions enables further improvements to predictive accuracy, as the LLM updates elicited distributions as we would expect. AutoElicit can therefore also be used to simplify the typical expert prior elicitation process by only requiring experts to provide insights in natural language. We find that prior elicitation with AutoElicit is an improved alternative to in-context learning in our setting, as we can inspect priors, confidently perform Bayesian inference, and often achieve reduced error.
On our private dataset, we find a large and quantifiable performance improvement when predicting cases of UTI. Within our study on dementia care, the resulting model created with LLM-elicited priors can provide better predictions of adverse health conditions, reliable uncertainty estimates, and remain interpretable and trusted by clinicians.
Impact Statement
This paper presents work that aims to make expert prior elicitation more accessible and suggests a new way for language models to be used in settings where high levels of trust in predictive models are required. We see many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@
AC: Original idea conception, software, analysis of results, writing, editing, reviewing, funding acquisition. RGK: Original idea conception, supervision, analysis of results, editing, reviewing, funding acquisition. PB: Supervision, analysis of results, reviewing, funding acquisition.
@@ Acknowledgments
This study is funded by the UK Dementia Research Institute (UKDRI) Care Research and Technology Centre funded by the Medical Research Council (MRC), Alzheimer’s Research UK, and Alzheimer’s Society (grant number: UKDRI–7002). Further funding was provided by the UKRI Engineering and Physical Sciences Research Council (EPSRC) PROTECT Project (grant number: EP/W031892/1) and the Royal Academy of Engineering, UK. Infrastructure support for this research was provided by the NIHR Imperial Biomedical Research Centre (BRC) and the UKRI Medical Research Council (MRC). Additionally, funding was provided by the Turing scheme, UK, and resources used in preparing this research were provided, in part, by the Province of Ontario, the Government of Canada through CIFAR, and companies sponsoring the Vector Institute. Moreover, RGK is supported by a Canada CIFAR AI Chair and a Canada Research Chair Tier II in Computational Medicine. The funders were not involved in the study design, data collection, data analysis, or manuscript writing.
@@ A.1 Availability of Code
We provide the code to generate all figures and results [^2] https://github.com/alexcapstick/llm-elicited-priors, and functions to use our method on new datasets [^3] https://github.com/alexcapstick/autoelicit, on GitHub.
This open-source code makes it simple to perform expert prior elicitation using LLMs: {minted}[mathescape, numbersep=5pt, frame=lines, framesep=5mm, ]python # import the elicitation function from llm_elicited_priors.gpt import ( get_llm_elicitation_for_dataset )
@@
This code is implemented in python (v3.11.9), using pip (v24.0). For our experiments, we use the following packages:
@@ A.2 Further Details on the Related Work
In this section, we outline in more detail the prior work in our area of research.
@@ A.2.1 Using LLMs to Elicit Expert Priors
Choi et al. (2022) explores the use of an LLM to perform feature selection, causal discovery, and to modify the reward function in a reinforcement learning task. In the first case, they observe that when listing features from two datasets, GPT-3 can correctly identify the features that correspond to a single specified predictive task. Additionally, when applied to causal discovery, the language model is often able to predict the directionality of causal relationships in the Tüebingen Cause-Effect Pairs dataset (Mooij et al., 2016). This inspires us to explore the possibility of using LLMs to elicit prior parameter distributions over linear models.
Gouk & Gao (2024) use LLMs to generate synthetic data to elicit priors on public datasets. A logistic regression model is then pre-trained using this data, defining a prior. However, this work does not discuss the memorisation of public datasets by LLMs, which is shown to be substantial (Bordt et al., 2024), supported by the results in Appendix A.8. In Appendix A.9.5, we compare the posterior performance of AutoElicit with the method proposed in Gouk & Gao (2024) and show that on private data, AutoElicit provides more informative priors.
Further, Selby et al. (2025) elicit beta distribution parameters to define a prior for predictive tasks where results from human experts are available. The priors from four LLMs were visually compared to those elicited from human experts using the SHELF (Oakley & O’Hagan, 2019; Gosling, 2018) method, demonstrating significant differences between experts and LLMs. However, by considering prior predictive distributions, they show that GPT-4 (OpenAI et al., 2024) can provide parameter distributions that allow for zero-shot performance comparable to models trained on tens of data points. These priors are over the targets directly, as with in-context learning, and are not updated using data. In our work, we use the LLM to elicit multiple Gaussian distributions, which are combined to create a mixture model for a more general distribution over the parameters of a linear predictive model. Our priors can then be updated using observed data. Further, we explore the collaboration of LLMs and human experts to provide more informative prior distributions.
In addition, Requeima et al. (2024) prompt language models with task descriptions and collected data points to generate posterior predictive distributions for regression tasks, which they find outperform Gaussian processes in some tasks. This is done by repeatedly sampling a language model at a new feature value, given the previous input and output observations, and a given text description that contains prior information from a human expert. In our work, we explore whether repeatedly sampling language models can also produce general prior distributions for predictive models.
@@
Other works explore using LLMs for defining entire modelling pipelines (Tornede et al., 2023; Zhang et al., 2023; Li et al., 2024). In particular, Li et al. (2024) questioned whether Box’s Loop (Box & Hunter, 1962) could be improved upon by using LLMs. In this work, they were used as model builders and critics to automate the model design and iteration steps. The authors find that this process can produce models that perform well in various modelling tasks and are comparable to those produced by domain experts. This demonstrates that LLMs have a good understanding of probabilistic modelling, suggesting that they might also be able to provide priors over predictive models directly.
Li et al. (2023) discusses using LLMs to produce prior distributions for perception and control tasks. This work focuses on vision and robotics and elicits common-sense priors from LLMs to bias perception and action models towards realistic world states. The authors find that for vision tasks, this improves predictions on rare, out-of-distribution, and structurally novel inputs. Our work aims to extend this by evaluating whether LLMs can also provide expert information about parameters in predictive models.
@@ A.2.2 In-context Learning
Learning from demonstrations in-context has proven to be effective for several natural language tasks, demonstrating the surprising ability of language models to perform outside of their intended use (Brown et al., 2020; Dong et al., 2024; Hegselmann et al., 2023; Chowdhery et al., 2023; Touvron et al., 2023). In Brown et al. (2020), the authors study GPT-3’s strong ability on natural language tasks after seeing a few examples, one example, or none at all. This work inspired others to study the extent to which language models can learn from demonstrations given as text alongside a task description. Further, Xie et al. (2021) study the mechanics of in-context learning and describes it from a Bayesian Inference perspective. They find that in-context learning succeeds when language models can infer a shared concept across individual examples in a prompt. Min et al. (2022) argue that in-context learning allows the language model to observe the label space, the distribution of input text, and the task format, as they find that swapping the real labels in demonstrations for random labels does not significantly reduce accuracy. Most of the existing work focuses on natural language tasks (Dong et al., 2024), leaving numerical problems less explored and a focus of our work.
In Hollmann et al. (2022), the authors train a transformer model to solve synthetically generated classification tasks from a prior over tabular datasets. Given a new task and demonstrations, this transformer performs in-context learning to make predictions using inference steps only. Furthermore, Akyürek et al. (2022) and Von Oswald et al. (2023) discuss the types of learning algorithms that language models (and more generally, transformers) can implement in-context. Focusing on linear models, Akyürek et al. (2022) show that, by construction, transformers are capable of performing gradient descent and closed-form ridge regression. Akyürek et al. (2022) also demonstrate that trained in-context learners closely match predictors trained through gradient descent. Further, Von Oswald et al. (2023) show that transformers, by incorporating multilayer perceptrons into their architectures, can also solve non-linear regression tasks, and Dalal & Misra (2024) propose a Bayesian framework to reason about a language model’s internal state and suggest that in-context learning is consistent with Bayesian learning. Wang et al. (2023), by viewing LLMs as latent variable models, suggest that in-context learning can reach the Bayes optimal predictor with a finite number of demonstrations, which can be chosen from a dataset using their framework. However, the authors Falck et al. (2024) provide evidence that LLMs violate the martingale property, which they show is a requirement of a Bayesian learning system for exchangeable data. Through this, they argue that the in-context learning behaviour LLMs exhibit is, in fact, not Bayesian.
In contrast to the literature, in our work we empirically test, through the approximation of the internal model used by a language model for its in-context predictions, whether for a given task the language model is performing Bayesian inference before and after seeing demonstrations.
@@ A.3 Formal Details on Extracting the Internal Prior and Posterior
In this section, we describe the method presented in Section 4.2 in more detail.
@@
A.4 Does the LLM use the Model Class we Specify?
As part of comparing AutoElicit with in-context learning, we use maximum likelihood sampling to extract the internal parameter distributions (discussed in Section 5.4), which we use to test whether the LLM is: (1) providing a similar prior distribution through elicitation to the one it is using in-context, and (2) empirically test whether the LLM is approximating Bayesian inference. However, when extracting the internal parameters of the in-context model, we assume that the LLM is applying the model class that we specify in the task description.
@@ A.4 Does the LLM use the Model Class we Specify?
In Figure 8 we show the true $\hat{y}_{n}$ against the in-context predictions $y_{n}$. Interestingly, even given the simple linear equation we ask the language model to mimic, we find that it struggles to provide good predictions in-context. This suggests that the language model is not always using the model class for in-context predictions that we specify in the task description, indicating that, in its current state, in-context learning might be inadequate for predictive tasks with primarily numerical features. This figure also shows that the language model was hesitant to predict negative values and preferred positive predictions – a strange behaviour that clearly restricts its use as an emulator of a linear model.
@@ A.7 Further Experimental Details
In all experiments, we use OpenAI’s GPT-3.5-turbo (specifically GPT-3.5-turbo-0125), GPT-4-turbo (GPT-4-turbo-2024-04-09) [^4] https://platform.openai.com/docs/api-reference/chat, and DeepSeek-R1-32B-Q4555https://ollama.com/library/deepseek-r1:32b with a temperature of $0.1$ and all other settings as default.
@@ A.7.1 Dataset Details
  • A private dataset collected as part of our study on Dementia care [^6] https://www.imperial.ac.uk/uk-dri-care-research-technology/: This dataset contains $10$ features engineered from measurements of in-home activity and physiology. The raw features were measured using passive infrared (PIR) sensors located around the homes of people living with dementia, and a sleep mat positioned in the participants’ beds. The targets correspond to days with a clinical labelling of positive or negative urinary tract infection (UTI) which were calculated through the analysis of a provided urine sample in consultation with a healthcare professional (Capstick et al., 2024). For this analysis, we used a subset of the data containing $78$ people and $371$ labelled days of UTI, with $270$ negatives and $101$ positives. When training and testing on this dataset, we centred and scaled the training data and used the same statistics to normalise the test data.
@@
  • The Diabetes dataset (Efron et al., 2004): After preprocessing, this dataset contains physiological measurements and demographic information from $442$ people with diabetes. Here, the task is to predict a quantitative measure of their diabetes disease progression one year after this baseline. This dataset consisted of $10$ features, and the targets were regression values with a maximum value of $10.0$, a minimum value of $0.0$, and a mean value of $3.96$.
@@
The UTI study was performed in collaboration with Imperial College London and Surrey and Borders Partnership NHS Trust. Participants were recruited from the following: (1) health and social care partners within the primary care network and community NHS trusts, (2) urgent and acute care services within the NHS, (3) social services who oversee sheltered and extra care sheltered housing schemes, (4) NHS Community Mental Health Teams for older adults (CMHT-OP), and (5) specialist memory services at Surrey and Borders Partnership NHS Foundation Trust. All participants provided written informed consent. Capacity to consent was assessed according to Good Clinical Practice, as detailed in the Research Governance Framework for Health and Social Care (Department of Health 2005) and the Mental Capacity Act 2005. Participants were provided with a Participant Information Sheet (PIS) that includes information on how the study uses their personal data collected in accordance with the GDPR requirements. If the participant was deemed to lack capacity, a personal or professional consultee was sought to provide written consent to the study. Additionally, the capacity of both the participant and study partner is assessed at each research visit. Research staff conducting the assessment have completed the NIHR GCP training and Valid Informed Consent training. If a participant is deemed to lack capacity but is willing to take part in the research, a personal consultee is sought in the first instance to sign a declaration of consent. If no personal consultee can be found, a professional consultee, such as a key worker, is sought. This process is included in the study protocol, and ethical panel approval is obtained.
@@ A.7.2 Language Model Details
These represent three distinct types of language models. GPT-3.5-turbo is a cost-effective option for those who are able to use OpenAI’s API and removes the need to run the language model locally. Similarly, GPT-4-turbo is hosted by OpenAI, however, it is a considerably more expensive option. In contrast, DeepSeek-R1-32B-Q4 is an open-source model that can be run locally, given the compute requirement is met, allowing priors to be elicited in a secure environment and without API costs.
@@ A.7.3 Evaluating LLM-elicited Priors
To evaluate the usefulness of the AutoElicit priors (presented in Section 4.1), we started by comparing their posterior performance to an uninformative prior. To do this, we chose the predictive tasks described in Section 5.1: $y=2x_{1}-x_{2}+x_{3}$, Urinary Tract Infection (UTI), Heart Disease, Hypothyroidism, Diabetes, and Breast Cancer and calculated the posterior accuracy or mean squared error on a test set after observing varied numbers of examples. For each of the datasets, we selected $50\%$ of the data points to form the test set and used the remaining data points to form a training set. This is done $10$ times to produce $10$ folds of random training and testing splits.
@@
To produce the results in Figure 9, we then sample the posterior predictive distribution over our test set and calculate the accuracy or mean squared error for each posterior parameter sample. To produce Figures 2 and 14, we also calculate the mean accuracy and the mean squared error for each of the test splits before plotting the mean and the $95\%$ confidence interval.
As an example, for the synthetic task, one of the task descriptions is as follows:
@@
  • User role: I am a data scientist working with a dataset for a prediction task using feature values to predict a target. I would like to apply your model to predict the target for my samples. My dataset consists of these features: [’feature 0’, ’feature 1’, ’feature 2’]. All the values have been standardised using z-scores. The known relationship is: ’target’ = 2 * ’feature 0’ - 1 * ’feature 1’ + 1 * ’feature 2’. Based on the correlation between each feature and target, whether positive or negative, please guess the mean and standard deviation for a normal distribution prior for each feature. I need this for creating a linear regression model to predict the target. Provide your response as a JSON object with the feature names as keys, each containing a nested dictionary for the mean and standard deviation. A positive mean suggests positive correlation with the outcome, negative for negative correlation, and a smaller standard deviation indicates higher confidence. Only respond with JSON.